# 3D-Prover: Diversity Driven Theorem Proving With Determinantal Point Processes

### v1

@@ 3.3.2 Ablation Study
Table 5: Percentage of unique environment responses per node in miniF2F-valid (Mean $\pm$ Standard Error). Unique defines either syntactically distinct error messages or responses including at least one previously unseen subgoal. $\theta$ controls the temperature of the model scores when calculating quality. We see that 3D-Prover gives a higher diversity of environment responses, increasing with $\theta$.
@@
8 85.3 $\pm$ 0.1 89.9 $\pm$ 0.1 90.1 $\pm$ 0.1 91.1 $\pm$ 0.1
16 77.5 $\pm$ 0.1 84.1 $\pm$ 0.1 84.9 $\pm$ 0.1 85.5 $\pm$ 0.1
32 72.3 $\pm$ 0.2 76.3 $\pm$ 0.2 76.9 $\pm$ 0.2 77.5 $\pm$ 0.2
Table 6: Percentage of successful tactics per node resulting in unique subgoal(s) over miniF2F-valid (Mean $\pm$ Standard Error). $\theta$ controls the temperature of the model scores in 3D-Prover when calculating quality. We observe 3D-Prover gives a higher number of unique subgoals per tactic, leading to a more diverse set of proof paths, with larger $\theta$ controlling this.
Diversity
To examine the diversity of a proof attempt, we consider two metrics. For the first, as in Table 5, we look at the percentage of unique environment responses to tactics executed per node, including responses with unique errors. As it is difficult to select tactics guaranteed to be successful (see Table 4), an exploratory policy should generate tactics which result in more varied outputs, so as to better explore the space. Table 6 quantifies the likelihood of a successful tactic resulting in a new proof path, where we restrict only to successful tactics to account for the discrepancy in success rate between approaches. This gives a direct metric measuring the effectiveness of tactics in providing distinct proof paths to explore. We first observe that the Random baseline leads to higher diversity, as expected. 3D-Prover outperforms both baselines, giving more diverse responses for both valid and invalid tactics. As intended, increasing the parameter $\theta$ results in further improvements to diversity under these metrics.
3D-Prover
$K$ Top-$K$ Random $\lambda_{\tau}=0.1$ $\lambda_{\tau}=1.0$
8 206 $\pm$ 0.8 198 $\pm$ 0.9 155 $\pm$ 0.5 136 $\pm$ 0.5
16 220 $\pm$ 0.8 218 $\pm$ 0.9 176 $\pm$ 0.6 152 $\pm$ 0.5
32 224 $\pm$ 0.8 215 $\pm$ 0.8 191 $\pm$ 0.7 181 $\pm$ 0.6
Table 7: Tactic execution time in milliseconds over miniF2F-valid proof attempts (Mean $\pm$ Standard Error). $\lambda_{\tau}$ controls the time weighting of the quality score in 3D-Prover. 3D-Prover selects faster tactics on average, with larger $\lambda_{\tau}$ magnifying this.
Execution Time
Table 7 shows the execution time for tactics over miniF2F-valid transitions. Again we see that 3D-Prover outperforms both baselines, with the improvement increasing with more filtering. Increasing the time weight $\lambda_{\tau}$ results in further reductions to the average execution time, demonstrating the accuracy of the predictions, and that they can directly result in faster tactics when filtering.
4 Conclusion
Future work
One might consider structured DPPs (Kulesza, 2012), which operate at the tree level to select diverse paths, rather than the node level, which selects diverse edges. Continual learning of the transition model is another avenue, where training on new data as it is generated could lead to more accurate assessments of diversity and quality. Our approach could also be combined with a separate search algorithm such as HTPS (Lample et al., 2022), rather than BestFS. Testing larger models would be a natural extension, for both the transition model and the underlying tactic generator. Our methodology may also be useful to enhance search in domains beyond formal proving, such as code generation or game playing.
Summary
We introduce 3D-Prover, a method to augment proof search by filtering candidate tactics to generate diverse and high quality subsets. By generating tactic representations which reflect the response of the proving environment, 3D-Prover is able to filter tactics based on their likely outcome. We evaluate 3D-Prover by augmenting the ReProver LLM on the standard miniF2F benchmark, where we find an improvement in the overall proof success rate (Table 2), particularly for deeper searches. Our ablation studies confirm the utility of our tactic representations, which allow the selection of tactics with improved success rates, diversity, and/or execution time. By effectively pruning the search space, 3D-Prover is a step towards enabling deeper automated proofs.
Appendix A Appendix
A.1 Pass@k
$K$ Random 3D-Prover Gain
8 25.7 28.6 +11.3%
16 30.2 31.0 +2.6%
32 29.8 29.8 +0.0%
Table 8: Percentage of proofs found after four attempts (Pass@4) on miniF2F-test, with $K$ tactics selected per node.
Table 8 shows the Pass@4 results for miniF2F-test, which is the number of proofs found at least once over four attempts. We compare 3D-Prover to the Random baseline, taking the same four runs from Table 2, where $\lambda_{s}=\lambda_{\tau}=0$, $\theta=1$. With Top-$K$ being deterministic, the Pass@$k$ rate is the same as the Pass@1 rate. Given several attempts, $K=16$ appears to provide a good tradeoff between breadth and depth, performing the best overall. 3D-Prover maintains a large improvement for $K=8$, with a modest improvement for $K=16$.
A.2 Proof success rate over hyperparameters
($\lambda_{s}$, $\lambda_{\tau}$, $\theta$)
8 25.0 25.0 25.8 22.5 23.8
16 29.1 28.7 27.9 27.0 26.6
32 28.7 28.3 28.7 27.9 27.0
Table 9: Pass@1 results on miniF2F-valid, over different hyperparameter configurations for 3D-Prover.
A.3 Embedding Discussion
Embedding Comparison
Embedding Objective
As outlined in Section 2, we train our embeddings to be reflective of the tactic semantics across all three components of Status, Time and Output. Hence 3D-Prover, which selects diverse embeddings, may lead to tactics predicted to have errors, where the errors are diverse in terms of their predicted message. The hyperparameter $\lambda_{s}$ can alleviate this by weighting the scores based on their likelihood of success. From our experiments (Table 9), there is not necessarily a benefit to Pass@1 by filtering out strongly based on the predicted error likelihood. To speculate, the error prediction, although quite good, is imperfect with many false negatives (Table 1). This can lead to potentially useful tactics being ignored if the error prediction is overly trusted, even though there is a higher tactic success rate overall as in Table 4. Given these prediction errors, it may be the case that selecting goals which are predicted to lead to (diverse) errors may be preferable, given the possibility they result in successful new subgoals. These subgoals may be be quite different from those previously selected, as they are mispredicted, so are clearly outside the space of tactics where the transition model is confident about the outcome. Further analysis could be worthwhile to investigate this. An embedding architecture trained only on successful tactics could be used, however given the high error rate of tactics, this would ignore a large proportion of the transition data.
A.4 Computational Overhead
On our hardware, we found 3D-Prover adds a constant overhead, taking approximately 2x as long for tactic generation. The majority of this is in generating embeddings for the 64 tactics, which we were unable to batch on our hardware due to memory constraints. The DPP algorithm itself added almost no overhead once the embeddings were generated. This could be sped up by batching (if memory permits), or through a different architecture. For example, the Separate model in Section 2.2 could be used, where tactics can be batched with much less memory. An augmented architecture which embeds the goal in isolation, which is then given to the tactic encoder as a single vector, could be used. This would provide a speed up while allowing some attention between the tactic and the goal, although not to the degree allowed for by our Combined model. As a proof of concept, we used the Combined model as it provides the most goal-aware embeddings to test our filtering algorithm.

### v2

@@ Appendix F Additional Diversity Analysis
Table 13: Average cosine similarity between subgoal embeddings for ReProver over miniF2F-valid (Mean $\pm$ Standard Error). We observe consistently lower similarity among subgoals generated from 3D-Prover, increasing with the diversity parameter $\theta$.
@@
8 0.939 $\pm$ 0.000 0.945 $\pm$ 0.000 0.931 $\pm$ 0.000 0.930 $\pm$ 0.000
16 0.916 $\pm$ 0.000 0.926 $\pm$ 0.000 0.910 $\pm$ 0.000 0.905 $\pm$ 0.000
32 0.904 $\pm$ 0.000 0.909 $\pm$ 0.001 0.900 $\pm$ 0.001 0.896 $\pm$ 0.001
Appendix G Computational Overhead
3D-Prover adds a constant but minimal time and memory overhead, the majority of which is in generating embeddings for the candidate tactics. Taking our first run for 3D-Prover with InternLM, we found the filtering time over nodes to be $0.07$s with a standard deviation of $0.02$s. In comparison, the average tactic generation time was $7.5$s with a standard deviation of $4$s. This gives us a time overhead of approximately $0.9\%$. The memory overhead was approximately 3GB of VRAM, while the tactic model took 44GB of VRAM, giving approximately 7% memory overhead.
We also note that the average time to execute a tactic from the No Filtering model was approximately $0.13$s (with $0.12$ standard deviation). With up to 64 candidate tactics per node, the filtering time of $0.07$ seconds is around half the average time to execute a single tactic. Given the improvements in success rate (Table 5), our filtering model therefore gives an effective way to reduce the total computational resources by preventing the wasteful execution of erroneous tactics. To further support this, we observed the average total time for a proof search attempt with 3D-Prover with InternLM ($K=8$) to be 993.8 seconds (1761.5 Standard Deviation), while the total time for Top-$K=8$ was 1116.8 seconds (2126.0 Standard Deviation).
Appendix H Number of tactic candidates for InternLM
As noted in 4(a), our experiments with InternLM2.5-Step-Prover use sampling rather than beam search for tactic generation, as done in Wu et al. (2024). As a result, there is no guarantee there will be 64 unique tactics, with many samples being identical. As a result, we sample 128 initial tactics, and take the total unique candidates as our ground set (up to 64 total). We plot the distribution of unique tactic candidates per node in Figure 6. This informed our choice of $K=8$ for our experiments, as higher $K$ would give minimal filtering in comparison to a beam search approach.
Figure 6: Distribution of unique tactic candidates per node for InternLM2.5-Step-Prover over miniF2F-test ($\mu=35.4,\sigma=11.5$).
Appendix I Computational Resources and Usage
For traning our transition models, we used a singe RTX4090 GPU with an Intel i9 13900k processor. For evaluating, we used two internal machines each with two RTXA6000 GPUs, and a Intel Xeon W-2223. For each evaluation experiment in 4(a), we assigned a single RTX A6000 GPU which contained both the tactic generator and transition model, which served 2 CPU provers which request tactics and evaluate in the Lean environment.
For each transition model, training for 2 epochs took approximately 2 days for each run, giving a total of 10 days of 4090 training for our results in 2.2. Each evaluation run for ReProver over miniF2F took around 12h, while each evaluation run for InternLM2.5-Step-Prover around 2 days. Counting the number of runs for all baselines and 3D-Prover, we have 28 runs for ReProver with miniF2F-test, 22 for miniF2F-valid and 17 runs for InternLM. The additional result over the large LeanDojo dataset in D took around 5 days per run, with 7 runs total. We therefore estimate the total evaluation time (for a single machine with 2 RTX6000s and a Xeon W-223) to be 94 days, which was around 47 days with our 2 machines. The full research project included additional compute and experiments, as different architectures and approaches were prototyped, however we do not have an estimate for the amount.